# Processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in Inert Higgs Doublet Models and Two Higgs Doublet Models

### v2

@@ 2.2 Two Higgs Doublet Models
[…]
3 One-loop corrections to $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in HESM
The calculations are performed with the help of FeynArts/FormCalc/LoopTools/Collider packages [102, 100, 101]. We first implement the above-mentioned HESMs into FeynArts [102] model. All one-loop diagrams for the computed processes are then generated automatically by using the program. In the frameworks of the HESM under investigations, all one-loop induced Feynman diagrams for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ are listed in the following paragraphs. The computations are performed in the ’t Hooft-Feynman gauge (HF), all Feynman diagrams can be categorized into several groups listed in next paragraphs. We first mention one-loop induced Feynman diagrams with $\phi_{k}^{*}$-poles for $\phi_{k}^{*}=h^{*},H^{*}$. These topologies are connected loop-induced processes $\phi_{k}^{*}\rightarrow\gamma\gamma$ with the vertices of $\phi_{k}^{*}\phi_{i}\phi_{j}$, as plotted in Fig. 1. In Fig. 1, all fermions exchanging in the loop are included in the group $G_{1}$. Feynman diagrams with vector $W$-boson propagating in the loop are concerned. Within the HF gauges, all charged Goldstone $G^{\pm}$ and Ghost particles exchanging in the loop are also taken into consideration this case. These diagrams are putted into the group $G_{2}$. We also have singly charged Higgs appear in the models under concern which they are also exchanged in the loop of diagrams, as noted as the group $G_{3}$.
Figure 1: One-loop induced Feynman diagrams with $\phi_{k}^{*}$-poles for $\phi_{k}^{*}=h^{*},H^{*}$.
One-loop box diagrams contributing to the computed processes are next discussed. Depend on the kind of particles propagating in the loop, we list the box diagrams into several groups as follows. In the group $G_{4}$, as plotted in Fig. 2, all fermions in the loop are concerned.
Figure 2: Box diagrams with fermions in the loop involving to the calculated processes in the HESM.
Additionally, conindering vector $W$-boson, charged Goldstone bosons and Ghost particles propagating in the loop of one-loop box diagrams, as plotted in Figs. 34, are also contributed to the processes under consideration. We classify these diagrams into the group $G_{5}$.
Figure 3: Vector $W$-boson, charged Goldstone bosons and Ghost particles propagating in the loop of one-loop box diagrams.
Figure 4: Vector $W$-boson, charged Goldstone bosons and Ghost particles propagating in the loop of one-loop box diagrams.
Within the frameworks of the HESMs under investigations in the paper, we also have one-loop box diagrams with mixing of vector $W$-boson and charged Higgs in the internal lines, seen Figs. 5,  6. We classify these diagrams into group $G_{6}$.
Figure 5: One-loop box diagrams with both vector $W$-boson, charged Goldstone bosons, Ghost particles and charged Higgs propagating in the loop.
Figure 6: Mixing of vector $W$-boson and charged Higgs in the internal lines of one-loop box diagrams.
Finally, we consider all one-loop box diagrams with singly charged Higgs in the loop, as shown in Fig. 7. These diagrams are then putted into group $G_{7}$.
Figure 7: One-loop box diagrams with charged Higgs in the loop propagating in the loop.
In general, one-loop amplitude for scattering processes $\gamma_{\mu}(q_{1})\,\gamma_{\nu}(q_{2})\rightarrow\phi_{i}(q_{3})\,\phi_{j}(q_{4})$ is expressed in terms of Lorentz structure as follows:
$\mathcal{A}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}$ $=$ $\Big{[}F_{00}\,g^{\mu\nu}+F_{12}\,q_{1}^{\nu}q_{2}^{\mu}+F_{13}\,q_{1}^{\nu}q_{3}^{\mu}+F_{23}\,q_{2}^{\mu}q_{3}^{\nu}+F_{33}\,q_{3}^{\mu}q_{3}^{\nu}\Big{]}\varepsilon_{\mu}(q_{1})\varepsilon_{\nu}(q_{2}).$ (28)
In this formulas, the vector $\varepsilon_{\mu}(q)$ is polarization vector of external photon with the $4$-dimension momentum $q$. The invariant masses are given: $q_{1}^{2}=q_{2}^{2}=0,\quad q_{3}^{2}=M_{\phi_{i}}^{2},\quad q_{4}^{2}=M_{\phi_{j}}^{2}$. The factors $F_{ij}$ for $i,j=1,2,3$ are called as one-loop form factors hereafter. They are presented as functions of the following kinematic invariant variables:
$s$ $=$ $(q_{1}+q_{2})^{2}=q_{1}^{2}+2q_{1}\cdot q_{2}+q_{2}^{2}=2q_{1}\cdot q_{2},$ (29)
$t$ $=$ $(q_{1}-q_{3})^{2}=q_{1}^{2}-2q_{1}\cdot q_{3}+q_{3}^{2}=M_{\phi_{i}}^{2}-2q_{1}\cdot q_{3},$ (30)
$u$ $=$ $(q_{2}-q_{3})^{2}=q_{2}^{2}-2q_{2}\cdot q_{3}+q_{3}^{2}=M_{\phi_{i}}^{2}-2q_{2}\cdot q_{3}.$ (31)
The kinematic variables obey the below relation as: $s+t+u=M_{\phi_{i}}^{2}+M_{\phi_{j}}^{2}$. Due to the on-shell photons in initial states, the amplitude follows the ward identity. Subsequently, the above-mentioned factors are related to each others as
$F_{00}$ $=$ $\dfrac{t-M_{\phi_{i}}^{2}}{2}\,F_{13}-\dfrac{s}{2}\,F_{12},$ (32)
$F_{00}$ $=$ $\dfrac{u-M_{\phi_{i}}^{2}}{2}\,F_{23}-\dfrac{s}{2}\,F_{12},$ (33)
$F_{13}$ $=$ $\dfrac{u-M_{\phi_{i}}^{2}}{s}\,F_{33},$ (34)
$F_{23}$ $=$ $\dfrac{t-M_{\phi_{i}}^{2}}{s}\,F_{33}.$ (35)
With the help of the mentioned relations, one-loop amplitude is written by two independent one-loop form factors, e.g $F_{12}$ and $F_{33}$ as chosen in this work. The one-loop amplitude can be finally presented in form of:
$\mathcal{A}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}$ $=$ $\Bigg{\{}\Bigg{[}\dfrac{(M_{\phi_{i}}^{2}-t)(M_{\phi_{i}}^{2}-u)}{2s}\cdot g^{\mu\nu}+q_{3}^{\mu}q_{3}^{\nu}+\dfrac{(t-M_{\phi_{i}}^{2})}{s}\cdot q_{2}^{\mu}q_{3}^{\nu}$
$\hskip 14.22636pt+\dfrac{(u-M_{\phi_{i}}^{2})}{s}\cdot q_{3}^{\mu}q_{1}^{\nu}\Bigg{]}\cdot F_{33}+\Big{[}q_{2}^{\mu}q_{1}^{\nu}-\dfrac{s}{2}\cdot g^{\mu\nu}\Big{]}\cdot F_{12}\Bigg{\}}\,\varepsilon_{\mu}(q_{1})\varepsilon_{\nu}(q_{2}).$
From the FeynArts program, we can generate the amplitude automatically which is expressed in term of one-loop tensor integrals. The mentioned tensor integrals appear in the production amplitude can be reduced into the scalar PV-functions by using the FormCalc package [103]. Finally, we collect all factors $F_{12}$ and $F_{33}$ presenting via the scalar PV-functions following the output of the packages LoopTools [100] and Collier [101].
The form factors $F_{ab}$ for $ab={12,33}$ are divided into several parts which are corresponding to the contributions from one-loop triangle and one-loop box diagrams given in the above-paragraphs. In detail, the form factors are decomposed as follows:
$F_{12}$ $=$ $\sum\limits_{\phi_{k}^{*}=h^{*},H^{*}}\dfrac{g_{\phi_{k}^{*}\phi_{i}\phi_{j}}}{\Big{[}s-M_{\phi_{k}}^{2}+i\Gamma_{\phi_{k}}M_{\phi_{k}}\Big{]}}\times$
×$\times\Bigg{\{}\sum\limits_{f}g_{\phi_{k}^{*}ff}\cdot F_{12,f}^{\text{Trig}}+g_{\phi_{k}^{*}WW}\cdot F_{12,W}^{\text{Trig}}+g_{\phi_{k}^{*}SS}\cdot F_{12,S}^{\text{Trig}}\Bigg{\}}$
$+\sum\limits_{f}g_{\phi_{i}ff}\cdot g_{\phi_{j}ff}\cdot F_{12,f}^{\text{Box}}$
+$+\Big{\{}g_{\phi_{i}WW}\cdot g_{\phi_{j}WW}\cdot F_{12,W}^{\text{Box, 1}}+g_{\phi_{i}\phi_{j}WW}\cdot F_{12,W}^{\text{Box, 2}}+g_{\phi_{i}\phi_{j}GG}\cdot F_{12,W}^{\text{Box, 3}}\Big{\}}$
+$+\Big{\{}g_{\phi_{i}SS}\cdot g_{\phi_{j}SS}\cdot F_{12,S}^{\text{Box},1}+g_{\phi_{i}\phi_{j}SS}\cdot F_{12,S}^{\text{Box},2}\Big{\}}$
$+g_{\phi_{i}SW}\cdot g_{\phi_{j}SW}\cdot F_{12,WS}^{\text{Box}},$
$F_{33}$ $=$ $\sum\limits_{f}g_{\phi_{i}ff}\cdot g_{\phi_{j}ff}\cdot F_{33,f}^{\text{Box}}+g_{\phi_{i}WW}\cdot g_{\phi_{j}WW}\cdot F_{33,W}^{\text{Box}}$
$+g_{\phi_{i}SS}\cdot g_{\phi_{j}SS}\cdot F_{33,S}^{\text{Box}}+g_{\phi_{i}SW}\cdot g_{\phi_{j}SW}\cdot F_{33,WS}^{\text{Box}},$
where $S\equiv H^{\pm}$ in this case. In further detail, the first part of form factors $F_{12}$ are calculated from one-loop diagrams of $\phi_{k}^{*}\rightarrow\gamma\gamma$ in connecting with the vertices $\phi_{k}^{*}\phi_{i}\phi_{j}$. These factors are presented in terms of each factor in the bracket, e.g. $F_{12,f}^{\text{Trig}}$ (contributed from the group $G_{1}$ in Fig. 1), $F_{12,W}^{\text{Trig}}$ (evaluated from $G_{2}$ in Fig. 1), $F_{12,S}^{\text{Trig}}$ (attributed from $G_{3}$ in Fig. 1). While the factors $F_{12,f}^{\text{Box}}$ are calculated from the fermion box diagrams in Fig. 2. Another form factors computed from one-loop $W$ box diagrams as in Fig. 3, are divided into the following parts, e.g. $F_{12,W}^{\text{Box, k}}$ for $k=1,2,3$ which are corresponding to the factors factorized out by general trilinear-couplings of $\phi_{i}WW,\leavevmode\nobreak\ \phi_{j}WW$, quadratic-couplings of $\phi_{i}\phi_{j}WW$ and $\phi_{i}\phi_{j}GG$ as in Eq. 3. We have also expressed the factors attributing from one-loop charged Higgs box diagrams into two sub-factors $F_{12,S}^{\text{Box},k}$ for $k=1,2$ which are factorized out by general trilinear-couplings of $\phi_{i}SS,\leavevmode\nobreak\ \phi_{j}SS$, quadratic-couplings of $\phi_{i}\phi_{j}WW$ and $\phi_{i}\phi_{j}SS$ as in Eq. 3. Lastly, from diagrams with mixing $W$ boson and charged Higgs in internal lines, we have the factors $F_{12,WS}^{\text{Box}}$ which can be factorized out in term of the trilinear-couplings $\phi_{i}SW,\leavevmode\nobreak\ \phi_{j}SW$. Otherwise, the form factors $F_{33}$ are only contributed from one-loop box diagrams. They can be factorized out in term of general couplings as in Eq. 3.
The factors $F_{12}$ and $F_{33}$ for the processes $\gamma\gamma\rightarrow hh$ in the HESM can be reduced to the results in the SM. For this case, we set all general couplings related to the HESM as $\phi_{k}SS,\leavevmode\nobreak\ \phi_{k}WS,\leavevmode\nobreak\ \phi_{i}\phi_{j}SS$ being zero and all other couplings is set back to the SM cases. Due to the $Z_{2}$-symmetry, the processes $\gamma\gamma\rightarrow hH$ in the IHDM are forbidden in this case. Reduction for factors $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM can be performed by setting appropriately the general couplings to the IHDM case. We can apply the same strategies for getting all one-loop form factors in the THDM. Having all the necessary factors for the computed processes, the tests for the calculations are performed, e.g. the ultraviolet finiteness, infrared finiteness of the one-loop amplitudes. Furthermore, the amplitudes also obey the so-called ward identity. This identity can be verified as follows. We collect all form factors $F_{00},\leavevmode\nobreak\ F_{12},\cdots$ independently and we then confirm the indentities from Eq. 32 to  35 numerically. Additionally, both the packages LoopTools and Collier are used for cross-checking for the final results. We skip showing the numerical results for the tests in this paper. For these topics, we refer our previous paper [129] for examples.
The cross-sections are then calculated as follows
[…]
with $n=2$ if the final particles are identical such as $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$, and 1 otherwise like $\gamma\gamma\rightarrow hH$. The integration limits are
$t_{\text{min}(\text{max})}$ $=$ $-\dfrac{s}{2}\Bigg{\{}1-\dfrac{M_{\phi_{i}}^{2}+M_{\phi_{j}}^{2}}{s}\pm\Bigg{[}1-2\Bigg{(}\dfrac{M_{\phi_{i}}^{2}+M_{\phi_{j}}^{2}}{s}\Bigg{)}+\Bigg{(}\dfrac{M_{\phi_{i}}^{2}-M_{\phi_{j}}^{2}}{s}\Bigg{)}^{2}\Bigg{]}^{1/2}\Bigg{\}}.$ (40)
The unpolarized amplitude is given
$\frac{1}{4}\sum\limits_{\text{unpol.}}\big{|}\mathcal{A}_{\gamma\gamma\rightarrow\phi_{i}\phi_{j}}\big{|}^{2}$ $=$ $\dfrac{M_{\phi_{i}}^{4}s^{2}+(M_{\phi_{i}}^{2}M_{\phi_{j}}^{2}-tu)^{2}}{8s^{2}}\,\big{|}F_{33}\big{|}^{2}$
$-\dfrac{M_{\phi_{i}}^{2}s}{4}\,\mathcal{R}e\Big{[}F_{33}\cdot\big{(}F_{12}\big{)}^{*}\Big{]}$
$+\dfrac{s^{2}}{8}\,\big{|}F_{12}\big{|}^{2}.$
@@
In phenomenological analyses, we are interested in examining the enhancement factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{NP}}$ with NP standing for the THDM and the IHDM, accordingly, defined as the ratio of cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESM over the corresponding ones for $\gamma\gamma\rightarrow hh$ in the SM. The factors are given explicitly by
@@
In this work, the enhancement factors are examined in the parameter space of the THDM and the IHDM.
4 Phenomenological results
For phenomenological results, all physical input parameters in the SM are given the same as in [128, 129]. Scanning parameters for each HESM will be shown in the next subsections.
@@ 4.1 IHDM
Phenomenological studies for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are presented in this subsection. In the IHDM, the process $\gamma\gamma\rightarrow hH$ is forbidden by the $Z_{2}$-symmetry. For this reason, we only concern physical results for the processes $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM.
@@ 4.1.1 Production cross-sections
In Fig. 8, we show cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM, together with the ones for $hh$ production in the SM, as functions of center-of-mass energy (CoM, or $\sqrt{\hat{s}_{\gamma\gamma}}$). For the generated data, we select the following parameter space in the IHDM as $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV and fix $\lambda_{2}=0.8$ for all cases. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV in the plots. Cross-sections are presented for $\mu_{2}^{2}=0$ GeV2 on the left panel and for $200^{2}$ GeV2 on the right panel, respectively. In the plots, the black line shows for cross-sections for $\gamma\gamma\rightarrow hh$ in the SM and the blue (green) line presents for $\gamma\gamma\rightarrow hh\leavevmode\nobreak\ (HH)$ in THDM, respectively.
@@
We first comment on the results in the case of $\mu_{2}^{2}=0$ GeV2. The cross-sections for $hh,\leavevmode\nobreak\ HH$ in the IHDM have peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. In the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV, $\hat{\sigma}_{hh,HH}$ in the IHDM are larger than $\hat{\sigma}_{hh}$ in the SM. Beyond the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, the cross-sections for $HH$ in the IHDM are suppressed in comparison with $hh$ productions in both the SM and the IHDM. It is interested in finding that the production cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM are dominant around the peaks compared with $\hat{\sigma}_{hh}$ in the SM. It indicates that the contributions from singly charged Higgs in the loop of $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are massive attributions in the regions.
@@
In the case of $\mu_{2}^{2}=200^{2}$ GeV2, we only observe a peak of $\hat{\sigma}_{HH}$ in the IHDM around $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. In further, the data shows that the cross-sections for $HH$ production are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV in contrast with the corresponding ones for $hh$ productions in both the SM and the IHDM. Beyond the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, $\hat{\sigma}_{HH}$ are suppressed. It is observed that the cross-sections for $hh$ production in the IHDM are smaller than $\hat{\sigma}_{hh}$ in the SM when $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV. In the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 550$ GeV, $\hat{\sigma}_{hh}$ in the IHDM tend to the cross-sections for $hh$ production in the SM. This can be explained as follows. Since the $hh$ productions in the IHDM are different from the ones in the SM by the contributions of charged Higgs in the loop of triangle $h^{*}$-pole and box diagrams. These contributions depend on $M_{H^{\pm}}$ and the vertices $hH^{\pm}H^{\mp}$, $hhH^{\pm}H^{\mp}$ expressing in terms of $\mu_{2}^{2}$. At the large value of $\mu_{2}^{2}$ these contributions may be cancelled out. As a result, cross-sections for $hh$ productions in the IHDM tend to the corresponding ones in the SM. Another case of $HH$ production, we have no couplings of $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ depend on $\lambda_{2}$. Therefore, we have no such cancellations as mentioned. It is reasonable that the cross-sections for $HH$ production in the IHDM are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV and they also have peak at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV.
@@
Figure 8: Total cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the SM and IHDM are presented as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$. For the generated data, we select $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV in this case. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. In the below plots, we fix $\mu_{2}^{2}=0,200^{2}$ GeV2 and $\lambda_{2}=0.8$ for all cases. In the plots, the black line shows for $\hat{\sigma}_{hh}$ in the SM. Additionally, the blue (green) line presents for $\gamma\gamma\rightarrow hh\leavevmode\nobreak\ (HH)$ in THDM.

### v3

@@ 2.2 Two Higgs Doublet Models
As in the case of the IHDM, we first summarize the current constraints on the parameter space of the THDM, given in Eq. 27. Both theoretical conditions and experimental data are taken into account, from which the allowed regions of the parameter space of the THDM are determined. Theoretical constraints arise from the requirements that the model remains within the perturbative regime, satisfies tree-level unitarity, and ensures vacuum stability of the scalar potential. These topics have been discussed in detail in Refs. [116, 117, 118, 119, 121] and references therein. We also consider electroweak precision tests (EWPT) in the context of the THDM, based on experimental data. Relevant implications for these constraints from LEP are reported in Refs. [122, 123]. The allowed mass ranges of the scalar particles in the THDM have been investigated at LEP, the Tevatron, and the LHC, as reviewed in Ref. [120]. Furthermore, one-loop induced decays of the SM-like Higgs boson, such as $h\rightarrow\gamma\gamma$ and $h\rightarrow Z\gamma$, have also been analyzed within the THDM in Refs. [112, 113] and references therein. By combining all of the above constraints, the physical parameter ranges can be chosen as follows: $126\leavevmode\nobreak\ \text{GeV}\leq M_{H}\leq 1000\leavevmode\nobreak\ \text{GeV}$, $60\leavevmode\nobreak\ \text{GeV}\leq M_{A^{0}}\leq 1000\leavevmode\nobreak\ \text{GeV}$, and $80\leavevmode\nobreak\ \text{GeV}\leq M_{H^{\pm}}\leq 1000\leavevmode\nobreak\ \text{GeV}$ for Type-I and Type-X THDMs. For Type-II and Type-Y, the physical parameters are scanned in the ranges: $500\leavevmode\nobreak\ \text{GeV}\leq M_{H}\leq 1000\leavevmode\nobreak\ \text{GeV}$, $500\leavevmode\nobreak\ \text{GeV}\leq M_{A^{0}}\leq 1000\leavevmode\nobreak\ \text{GeV}$, and $580\leavevmode\nobreak\ \text{GeV}\leq M_{H^{\pm}}\leq 1500\leavevmode\nobreak\ \text{GeV}$. The soft $Z_{2}$-breaking parameter is chosen as $m_{12}^{2}=M_{H}^{2}s_{\beta}c_{\beta}$. Finally, additional constraints in the $(t_{\beta},M_{H^{\pm}})$ plane are examined by incorporating flavor physics data, as shown in Ref. [124]. The results in Ref. [124] indicate that small values of $t_{\beta}$ are favored to be consistent with flavor experimental constraints. In our complementary analysis, we also explore the phenomenology associated with low $t_{\beta}$ scenarios.
3 Phenomenological results
@@
The detailed calculations for $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ with CP-even Higgses $\phi_{i,j}\equiv h,H$ in the HESM have been reported in our previous work [100]. In this paper, we focus on the phenomenological analysis of the concerned processes in the IHDM and the THDM. We are interested in examining the enhancement factors $\mu_{\phi_{i}\phi_{j}}^{\textrm{NP}}$ with NP standing for the THDM and the IHDM, respectively, defined as the ratio of the cross-sections of $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the HESM to the corresponding ones for $\gamma\gamma\rightarrow hh$ in the SM. The factors are given explicitly by
@@
In this work, the enhancement factors are examined within the parameter space of the THDM and the IHDM. Forthe phenomenological results, all physical input parameters in the SM are taken to be the same as those in [125, 126].
@@ 3.1 IHDM
Phenomenological studies for the processes $\gamma\gamma\rightarrow\phi_{i}\phi_{j}$ in the IHDM are presented in this subsection. In the IHDM, the process $\gamma\gamma\rightarrow hH$ is forbidden by the $Z_{2}$-symmetry. For this reason, we only focus on physical results for the processes $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM. We note that all physical results in the IHDM presented in the following subsections can be considered as the first outcomes from this work.
@@ 3.1.1 Production cross-sections
In Fig. 1, we show cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM, together with those for $hh$ production in the SM, as functions of center-of-mass energy (CoM, or $\sqrt{\hat{s}_{\gamma\gamma}}$). For the generated data, we select the following parameter space in the IHDM: $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV, and fix $\lambda_{2}=0.8$ for all cases. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV in the plots. Cross-sections are presented for $\mu_{2}^{2}=0$ GeV2 on the left panel and for $200^{2}$ GeV2 on the right panel, respectively. In the plots, the black line represents the cross-sections for $\gamma\gamma\rightarrow hh$ in the SM, and the blue (green) line represents the cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ (HH)$ in the IHDM, respectively.
@@
We first comment on the results in the case of $\mu_{2}^{2}=0$ GeV2. The cross-sections for $hh,\leavevmode\nobreak\ HH$ in the IHDM have peaks at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. In the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV, $\hat{\sigma}_{hh,HH}$ in the IHDM are larger than $\hat{\sigma}_{hh}$ in the SM. Beyond the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, the cross-sections for $HH$ in the IHDM are suppressed in comparison with $hh$ production in both the SM and the IHDM. It is interesting to find that the production cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the IHDM are dominant around the peaks compared with $\hat{\sigma}_{hh}$ in the SM. This indicates that the contributions from singly charged Higgs in the loop of $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ are significant in these regions.
@@
In the case of $\mu_{2}^{2}=200^{2}$ GeV2, we only observe a peak of $\hat{\sigma}_{HH}$ in the IHDM around $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV. Furthermore, the data shows that the cross-sections for $HH$ production are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 750$ GeV, in contrast with the corresponding ones for $hh$ production in both the SM and the IHDM. Beyond the regions $\sqrt{\hat{s}_{\gamma\gamma}}\geq 750$ GeV, $\hat{\sigma}_{HH}$ is suppressed. It is observed that the cross-sections for $hh$ production in the IHDM are smaller than $\hat{\sigma}_{hh}$ in the SM when $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV. In the regions of $\sqrt{\hat{s}_{\gamma\gamma}}\geq 550$ GeV, $\hat{\sigma}_{hh}$ in the IHDM tends to the cross-sections for $hh$ production in the SM. This can be explained as follows. Since the $hh$ production in the IHDM are different from those in the SM by the contributions of charged Higgs in the loop of triangle $h^{*}$-pole and box diagrams. These contributions depend on $M_{H^{\pm}}$ and the vertices $hH^{\pm}H^{\mp}$, $hhH^{\pm}H^{\mp}$, expressed in terms of $\mu_{2}^{2}$. At large values of $\mu_{2}^{2}$, these contributions may be cancelled out. As a result, cross-sections for $hh$ production in the IHDM tend to the corresponding ones in the SM. In another case of $HH$ production, we have no couplings of $HH^{\pm}H^{\mp}$ due to the $Z_{2}$-symmetry and the vertex $HHH^{\pm}H^{\mp}$ depends on $\lambda_{2}$. Therefore, we have no such cancellations as mentioned. It is reasonable that the cross-sections for $HH$ production in the IHDM are dominant in the regions $\sqrt{\hat{s}_{\gamma\gamma}}\leq 550$ GeV and they also have a peak at $\sqrt{\hat{s}_{\gamma\gamma}}\sim 2M_{H^{\pm}}=400$ GeV.
@@
Figure 1: Total cross-sections for $\gamma\gamma\rightarrow hh,\leavevmode\nobreak\ HH$ in the SM and IHDM are presented as functions of $\sqrt{\hat{s}_{\gamma\gamma}}$. For the generated data, we select $M_{H^{\pm}}=200$ GeV, $M_{H}=150$ GeV in this case. We vary $350$ GeV $\leq\sqrt{\hat{s}_{\gamma\gamma}}\leq 1500$ GeV. In the plots below, we fix $\mu_{2}^{2}=0,200^{2}$ GeV2 and $\lambda_{2}=0.8$ for all cases. In the plots, the black line shows $\hat{\sigma}_{hh}$ in the SM. Additionally, the blue (green) line presents $\gamma\gamma\rightarrow hh\leavevmode\nobreak\ (HH)$ in the THDM.